T2-5 Calculate equivalent units (LO 3) Sury Company reported the following information for the month of April:

Units

Work in Process
 Beginning inventory, April 1

 8,000

 Added to production

45,000

 Completed and transferred out

?

 Ending inventory, April 30

 5,000

Ending inventory was 100% complete with respect to materials and 80% complete with respect to conversion costs.

Required

a.How many units were transferred out of Work in Process Inventory?

b.What were the equivalent units of production for materials?

c.What were the equivalent units of production for conversion?

T2-6 Assign costs to completed units and ending Work in Process Inventory (LO 4) Refer to the information in T2-5. Sury also reported the following cost information in April:

Costs

Materials

Conversion

Work in Process
 Beginning inventory, April 1

$11,000

$ 17,160

 Added to production

 66,380

 104,000

Required

a.What was the cost per equivalent unit for materials?

b.What was the cost per equivalent unit for conversion?

c.What cost would be assigned to the units transferred out of Work in Process Inventory?

d.What cost would be assigned to the ending Work in Process Inventory?
Answer #1

1) Unit Completed and transferred out = 8000+45000-5000 = 48000

2) Equivalent unit of material = 48000+5000 = 53000

3) Equivalent unit of conversion = 48000+(5000*80%) = 52000

4) Cost per equivalent unit of material = (11000+66380)/53000 = 1.46

5) Cost per equivalent unit of conversion = (17160+104000)/52000 = 2.33

6) Cost of unit transferred out= 3.79*48000 = 181920

7) Cost of ending WIP = (5000*1.46+4000*2.33) = 16620